# Trajectory Encryption Cooperative Salvo Guidance

### v1

@@ I Introduction
In this paper, we consider a cooperative simultaneous interception of a fixed-wing drone leveraging fixed-wing and multirotor vehicles. The fixed-wing vehicles in the swarm have the ability to modulate their bearing angles to a constant value for time-constrained target interception, thus utilizing deviated pursuit. The multirotors, on the other hand, are endowed with the capability to modulate their thrust or propulsion, thus utilizing both radial and tangential components of their accelerations following true proportional navigation. Although different classes of vehicles have different time-to-go formulations, we show that consensus in their heterogeneous time-to-go values is established within a predefined time via the proposed method, regardless of the initial engagement geometry.
@@ II Preliminaries and Problem Formulation
where $t_{e}>0$ denotes a user-specified convergence time. The corresponding time derivative is given by
@@ III Trajectory Encryption Guidance Strategy
$a_{P_{k}}=\dfrac{-F_{k}+\left(-\alpha+\beta\dfrac{\dot{h}(t)}{h(t)}\right)\displaystyle\sum_{j=1}^{N}[\mathcal{L}]_{kj}e_{j}}{B_{k}},$ (11)
@@
From (15), we can observe that the system is globally asymptotically stable for $t\in(0,\infty)$ and from (17) it can be inferred that $V(t)\equiv 0$, which implies that $e(t)\equiv 0$ as $t\rightarrow t_{e}^{-}$. Therefore, from Lemma 1 we can conclude that the time-to-go error dynamics (10) is globally prescribed time stable under the proposed command $a_{P_{k}}$. This essentially means that a consensus is established in the pursuers’ time-to-go values within a time $t_{e}$ regardless of the initial engagement configuration. Thereafter, since $e_{k}$ is decreasing to zero, so is $t_{go_{k}}$, leading to a simultaneous interception. ∎
@@ IV Simulations
To further demonstrate the applicability of the proposed strategy, we employ the Predicted Interception Point (PIP)–based guidance strategy proposed in [7], in conjunction with DPG and TPNG. Consider the same engagement conditions as in the previous case, but pursuers $P_{1}$ and $P_{4}$ switch to the PIP-based guidance strategy, $P_{2}$ employs DPG, and $P_{3}$ does not switch and follows TPNG. From Figure 6(a), it can be observed that the pursuers intercept the target cooperatively at $58\text{s}$. Compared to the previous case, this demonstrates that cooperative interception remains feasible even when three different guidance laws are followed. Similar behaviors in the time-to-go and acceleration demands can be observed from Figure 6(b) and Figure 6(c), respectively, consistent with the previous case. Pursuers $P_{2}$ and $P_{3}$ exhibit velocity and lead angle characteristics consistent with the guidance law being followed, as shown in Figure 6(d). In contrast, for $P_{1}$ and $P_{4}$, the velocity remains constant after switching, while the lead angle ultimately converges to zero, which is consistent with [7].

### v2

@@ I Introduction
In this paper, we consider a cooperative simultaneous interception of a fixed-wing drone using fixed-wing and multirotor vehicles. However, any set of heterogeneous vehicles can be accommodated using the proposed strategy. The fixed-wing vehicles in the swarm have the ability to modulate their bearing angles to a constant value for time-constrained target interception, thus utilizing deviated pursuit. The multirotors, on the other hand, are endowed with the capability to modulate their thrust or propulsion, thus utilizing both radial and tangential components of their accelerations following true proportional navigation. Although different classes of vehicles have different time-to-go formulations, we show that consensus in their heterogeneous time-to-go values is established within a predefined time via the proposed method, regardless of the initial engagement geometry.
@@ II Preliminaries and Problem Formulation
where $t_{e}>0$ denotes a user-specified convergence time and $h(t)<0$ for $t\in[0,t_{e})$. The corresponding time derivative is
@@ III Trajectory Encryption Guidance Strategy
$a_{P_{k}}=\dfrac{{-1}-F_{k}+\left(-\alpha+\beta\dfrac{\dot{h}(t)}{h(t)}\right)\displaystyle\sum_{j=1}^{N}[\mathcal{L}]_{kj}e_{j}}{B_{k}},$ (11)
@@
From (15), we can observe that the system is globally asymptotically stable for $t\in(0,\infty)$ and from (17) it can be inferred that $V(t)\equiv 0$, which implies that $e(t)\equiv 0$ as $t\rightarrow t_{e}$. Therefore, from Lemma 1 we can conclude that the time-to-go error dynamics (10) is globally prescribed time stable under the proposed command $a_{P_{k}}$. This essentially means that a consensus is established in the pursuers’ time-to-go values within a time $t_{e}$ regardless of the initial engagement configuration. Thereafter, since $e_{k}$ is decreasing to zero, so is $t_{go_{k}}$, leading to a simultaneous interception. ∎
@@ IV Simulations
The proposed strategy is further validated without assuming the vehicle is a multirotor, we employ the Predicted Interception Point (PIP)–based guidance strategy proposed in [7] (requires control solely through lateral acceleration channel), in conjunction with DPG and TPNG. Consider the same engagement conditions as in the previous case, but pursuers $P_{1}$ and $P_{4}$ switch to the PIP-based guidance strategy, $P_{2}$ employs DPG, and $P_{3}$ does not switch and follows TPNG. From Figure 6(a), it can be observed that the pursuers intercept the target cooperatively at $58\text{s}$. Compared to the previous case, this demonstrates that cooperative interception remains feasible even when three different guidance laws are followed. Similar behaviors in the time-to-go and acceleration demands can be observed from Figure 6(b) and Figure 6(c), respectively, consistent with the previous case. Pursuers $P_{2}$ and $P_{3}$ exhibit velocity and lead angle characteristics consistent with the guidance law being followed, as shown in Figure 6(d). In contrast, for $P_{1}$ and $P_{4}$, the velocity remains constant after switching, while the lead angle ultimately converges to zero, which is consistent with [7].